OpenStudy (anonymous):

The cost per minute of a conference phone call, originally $2.80, increased by 20%. What is the new price? MEdal for correct answer!

OpenStudy (nurali):

10% of $2.80 is $.28 20% is double that, .$56 .$2.80 + .56 $3.36

OpenStudy (anonymous):

k can you answer some more

OpenStudy (nurali):

i will try

OpenStudy (anonymous):

$3.20, increased by 40%

OpenStudy (anonymous):

Okay so 40% of 3.20 is 1.28 so 3.20+1.28=4.48

OpenStudy (anonymous):

can you answer a few more

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Elena buys several shares of stock at $28 per share and sells them at $24.50 per share. What is the percent of decrease in the price of the stock?

OpenStudy (anonymous):

A. 11.5% B. 13.0% C. 12.0% D. 12.5%

OpenStudy (anonymous):

one second...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The stock suffered a 12.5% decrease

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

next may be a lil tricky

OpenStudy (anonymous):

Jack earned $972 in 6 years on an investment at a 3% annual simple interest rate. How much was Jack’s investment? A. $174.96 B. $486 C. $1944 D. $5400

OpenStudy (anonymous):

$5400

OpenStudy (anonymous):

$972/6 = $162/yr $162/.03 = $5400

OpenStudy (anonymous):

ok last one

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

A $6000 investment earns $690 annual simple interest in one year. What is the annual interest rate?

OpenStudy (anonymous):

what are your answers? Is it asking for the percentage?

OpenStudy (anonymous):

A. 10.5% B. 11% C. 11.5% D. 12%

OpenStudy (anonymous):

okay just one second...

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

11.5%

OpenStudy (anonymous):

Thanks You earned a medal + fan

OpenStudy (anonymous):

Awesome! Thanks a lot! :)

OpenStudy (anonymous):

got to go now cya!

OpenStudy (anonymous):

K keep studying hard!